OpenStudy (anonymous):

The distance traveled by an object can be modeled by the equation d = ut + 0.5at^2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work. I have the first steps down, but I'm completely confused on the other. The first step is to subtract both sides by ut, which gives me d-ut= 0.5at^2. The second step is to divide both side by 0.5, which then gives me d-ut over 0.5= at^2. But a still isn't by itself. Help??

OpenStudy (anonymous):

then divide both sides by t^2

OpenStudy (anonymous):

Oh, so it would be d-ut over .5t^2= a?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Ah. Thank you! I had a feeling it was that, but I didn't quite understand it

OpenStudy (campbell_st):

to get rid of 0.5 multiply both sides of your equation by 2 2(d - ut) = at^2 now divide

OpenStudy (anonymous):

no problem